A poil must be kept between 7.2 to 7.8. If a pool held 3.5 x 10^4L of water. Calculate
a) the minimum acceptable hydrogen ion concentration

b) the number of hydroxonium ions that must be present in the swimming pool when the pH is 7.8

c) the number of moles of HCl that would need to be added to the pool to change the pH from 7.8 to 7.2

d) the number of concentrated HCl ( 12mol L^-1) that would be need to be added to the pool to achieve this pH.
1
Expert's answer
2020-03-31T12:31:02-0400

a) the minimum acceptable concentration of [H+] refers to maximum acceptable pH = 7.8. From definition of pH it follows:

[H+] = 10-7.8 = 1.58*10-8 mol/L

b) the number of hydroxonium ions in the pool is obtained from the above concentration multiplied by volume

n = 1.58*10-8*3.5*104 = 5.53*10-4 mol, or multiplying by Avogadro number:

N = 5.53*10-4*6.02*1023 = 3.33*1018 ions

c) at pH = 7.2 there are n = 10-7.2*3.5*104 = 22.08*10-4 mol of H+

to reach this amount starting from the above case, one should add:

22.08*10-4 - 5.53*10-4 = 16.55*10-4 = 1.655*10-3 mol of HCl

d) dividing the above moles of HCl by its concentration we obtain that:

1.655*10-3/12 = 1.379*10-4 L of concentrated HCl should be added